OpenStudy (anonymous):

Simplify: 4x -12 divided by x^2 - 9 This is the end of my current section

OpenStudy (anonymous):

\[\frac{ 4x-12 }{ x^2-9 }\] \[\frac{ 4(x-3) }{ (x-3)(x+3) }\]

OpenStudy (anonymous):

can you take it from here or want me to elaborate more?

OpenStudy (anonymous):

i need to see the next step too

OpenStudy (anonymous):

the (x-3) is the common value on both the numerator and denominator so this leaves us with \[\frac{ 4 }{ (x+3)}\]

OpenStudy (anonymous):

you can cancel the (x-3)

OpenStudy (anonymous):

4/x+3....that was my numbers toobut wasn't for sure.... thanks

OpenStudy (anonymous):

no problem! Anytime